# Fast dynamical similarity analysis

### v1

@@ 2.3 Assessing novel optimizers for dynamic similarity analysis
We assessed the performance of the tested methods by computing four summary measures and showed that all were superior to DSA, with some variability among them. We included the following summary measures: (1) the alignment loss, $\|A_{1}-CA_{2}C^{\top}\|_{F}$ (Euclidean score), (2) an angular discrepancy between $A_{1}$ and the aligned $CA_{2}C^{\top}$ to capture qualitative misalignment, (3) the orthogonality residual $\|C^{\top}C-I\|_{F}$, and (4) wall–clock runtime. All hyperparameters, initializations, and stopping criteria were held fixed across methods. The runtime analysis indicated that the regularization-based variant and the Landing algorithm were the most computationally efficient across varying matrix sizes, with the regularization variant being marginally faster (Figure 3a). The Riemannian method was faster than the original DSA optimizer for smaller matrices ($n\approx 32$) but exhibited a steeper runtime growth, eventually surpassing DSA’s computational cost for larger values of $n$.
@@
Taken together, our analyses support the Landing–based optimization for solving the alignment problem as the preferred choice in most cases. It matches or improves upon the alignment accuracy of DSA, consistently outperforms both DSA and the penalty-based and Riemannian variants in wall–clock time, and maintains robust behavior across a wide range of matrix sizes.
2.4 Assessing efficacy and reliability of fastDSA
To demonstrate that the family of fastDSA methods is both efficient and reliable for dynamic similarity analysis we compare them under diverse scenarios. We apply them to two major sets of simulations [similar to ostrow2023beyond], where the underlying dynamics are well characterized a priori (and therefore can serve as a ground truth), and investigate (in)sensitivity, convergence, computational efficacy, and accuracy of all fastDSA variants to different forms of changes in governing dynamics.
2.4.1 Invariance under geometric deformations
We first investigated the insensitivity of fastDSA variants to purely geometrical transformations (i.e., those that leave the underlying dynamics intact). More specifically, we investigate whether fastDSA maintains invariance under geometric deformations that preserve the attractor topology.
The application of all fastDSA methods on geometric deformations that preserve the attractor structure suggests that fastDSA, similar to the original DSA [ostrow2023beyond], maintains invariance under these transformations. We compared the performance and computational efficiency of fastDSA with the original DSA based on simulation of a neural system known to represent head direction in the brain, which implements a ring attractor topology embedded in a network of $n$ neurons [wang2022multiple]. We use a conventional network architecture and drive the network using constant input and dynamic noise. Dynamics of the network result in an activity bump moving around the ring attractor at a roughly constant velocity (Figure 4a, left):
@@ 2.4.1 Invariance under geometric deformations
Our results suggest that all members of the fastDSA family perform equally well as the original DSA, and it is computationally more efficient. Comparing different variants of fastDSA and the original DSA demonstrates that all methods remain invariant across geometric deformations (Figure 4a, middle; red, green, and purple lines). In contrast, the Procrustes method rapidly diverges with higher values of $\beta$ (Figure 4a, middle; yellow). Furthermore, our results suggest, on average, fastDSAs computations are at least an order of magnitude faster ($\sim 15\times$) than the original DSA (Figure 4a, right).
The optimization behavior of the fastDSA family appears to mirror that of DSA, suggesting comparable optimization landscapes. To evaluate the convergence behavior of different fastDSA methods in comparison to DSA, we conducted a series of numerical experiments. We selected two distinct deformation magnitudes, $\beta=1$ and $\beta=4$ (from Figure 4a, middle). In both cases, we anticipated low dynamical similarity distances, reflecting minor or no topological changes. All fastDSA methods show a decreasing loss (Figure 4b, c left) and lower similarity scores (Figure 4b, c middle), reflecting insensitivity to geometrical deformation. Results indicated that fastDSA methods consistently converged to stable similarity scores, slightly slower than DSA.
Considering convergence and iteration time of fastDSA, the overall computational efficiency of fastDSA still remains superior to DSA. To more precisely characterize convergence speed, we analyzed when each algorithm’s loss curve flattened by collapsing run-by-run data into a single mean-loss trajectory per method. For each trajectory, we computed the discrete gradient as the absolute difference between successive loss values, $|\Delta\text{loss}|$, and used a small threshold (0.005) to define when changes become negligible. The stopping iteration was then defined as the first point where $|\Delta\text{loss}|$ dropped below this threshold, or alternatively, the iteration with the minimum gradient if no threshold was crossed. Despite converging slightly later, fastDSA variants was faster overall, requiring substantially less computational time per iteration compared to DSA (see, Figure 4b-c, right, for each iteration time).
@@ 2.4.2 Ability to detect shifts in governing dynamics
On the second step of evaluating fastDSA efficiency and reliability, we investigate the sensitivity of all fastDSA methods to changes in system dynamics; in particular, parametric transformations that alter the topology of the attractors in the dynamical system. To probe the sensitivity of fastDSA [similar to original DSA, ostrow2023beyond], we applied a parametrically controlled deformation of the ring topology into a line attractor within the same network model (Figure 4d, left). We implement it by progressively ablating neurons at one end of a recurrent neural network governed by a cosine connectivity kernel. We defined a continuous deformation parameter:
@@
Our results demonstrate that all suggested variants are equally sensitive to change in dynamics, as the original DSA, and it is computationally more efficient. Both the family of fastDSA and DSA sharply increased their estimated distance value as $\alpha$ transitioned from less than one (line attractor) to exactly one (ring attractor), confirming their sensitivity to topological changes (Figure 4d, middle; blue, red, green, and purple traces), in contrast to Procrustes analysis (yellow trace). Furthermore, comparing the computation times suggests that fastDSA was approximately 15 times faster than DSA in detecting these topological changes (Figure 4d, right).
@@ 2.4.3 Sensitivity to fine changes in governing dynamics
Lastly, we investigate whether fastDSA is sensitive to fine changes in dynamics and whether this sensitivity is superior to previously suggested alternatives [e.g., Wasserstein-based dynamics similarity, kamiya2024koopman, redman2024identifying]. A key advantage of fastDSA compared to alternative routes [in terms of computational efficiency, redmanconjugacy, redmanneurips, kamiya2024koopman] is its sensitivity to minor changes in the dynamics of the system, while maintaining the robustness to noise. The capacity of fastDSA to maintain sensitivity, as elaborated below, stems from its multifaceted approach to computing dynamic similarity. The FastDSA family and the DSA take into account both eigenvalues and eigenvectors to compare the governing dynamics; whereas Wasserstein-based approaches [redmanconjugacy, redmanneurips, kamiya2024koopman] focus only on eigenvalue distributions. Thus, as long as the eigenvectors shaping the flow fields are sufficiently structured and distinct, even in the presence of noise, the governing dynamics can be discriminated.
We first introduce one of the key routes for accelerating the estimation of dynamic similarity [discussed in, kamiya2024koopman], which is based on a combination of KernelDMD [Williams_2015, baddoo2022kernel], and using the Wasserstein distance as the similarity measure (we refer to it by kwDSA in the following), which brings more computational efficiency to dynamic similarity analysis. We then demonstrate that this alternative approach, kwDSA, is nevertheless less sensitive to changes in dynamics compared to all variants of fastDSA we introduced. Employing kernel-based methods [bernhard2004, bernhard2001] and KernelDMD [gonzalez2021kernel], more specifically, using kernel functions (e.g., Radial Basis Function or RBF) to implicitly map data into a high-dimensional feature space (known as reproducing kernel Hilbert space), considerably enhances computational efficiency, particularly for large-scale or high-dimensional data. This flexibility stems from representing complex nonlinearities without explicitly mapping to the high-dimensional space. By combining KernelDMD with the Wasserstein distance, the similarity between systems can be assessed by comparing the distributions of their spectral components, typically the eigenvalues or singular values of the corresponding Koopman operators [kamiya2024koopman]. We calculated KernelDMD as
@@
The Wasserstein distance provides a principled way to compare probability distributions and has been used to quantify spectral differences between dynamical systems. A recent study [redman2024identifying] demonstrated that Wasserstein-based metrics can capture dynamical similarity across nonlinear systems, showing both stability and interpretability when comparing spectral measures derived from Koopman or DMD operators. This suggest the Wasserstein distance could be a powerful tool for distribution-based comparisons. However, as we elaborated below, focusing solely on eigenvalue distributions can obscure important geometric differences encoded in the eigenvectors and the underlying flow fields. Thus, it possibly makes the dynamic similarity estimation less sensitive to differences that lie in the configurations of eigenvectors and the associated vector fields. Notably, the distribution of eigenvalues might be very similar across systems, even when the corresponding eigenvectors (which encode the flow directions) differ. To demonstrate the potential shortcomings of kwDSA, we compared the results of kwDSA with fastDSA variants.
@@
Our results suggest that kwDSA is less sensitive to the transition in the regime of the system’s dynamics as the parameter $\alpha$ approaches one (Figure 5a). The advantage of fastDSA over kwDSA is even more pronounced in the presence of noise in the data. To evaluate robustness to noise, we repeated the ring-to-line attractor transition experiment (presented in Figure 4d) while adding Gaussian white noise to the input data (observation noise). We tested four different signal-to-noise ratios (SNRs): 855, 213, 8.3, and 0.6 (corresponding to progressively increasing levels of noise). More specifically, zero-mean white noise was added directly to the neural activity trajectories before applying DMD-based computations. Across all noise levels, fastDSA consistently detected the topological transition (as $\alpha$ approached 1), while kwDSA was less sensitive (Figure 5a; detection capacity degraded after SNR 8.3, second column). These results further highlight fastDSA’s robustness to input variability and its advantage in capturing dynamical changes, even under noisy conditions, compared to fast alternatives, such as kwDSA. Although kwDSA can serve as a reasonable approximation in certain cases [for instance, where comparing only the eigenvalue spectra is sufficient, and we provided an improved implementation of kwDSA in our fastDSA package, behradcmclab], it might not be sufficiently sensitive to fine changes in the system’s dynamics; for instance (as discussed in the following) when the dynamics of the systems under comparison only have fine differences in their flow fields.
@@
Lastly, to explicitly demonstrate the mentioned limitation of kwDSA for dynamic similarity estimation (insensitivity to differences in eigenvectors), we also took an analytical approach. We defined two low-dimensional nonlinear systems (system A and system B) with identical eigenvalues but distinct eigenvectors (Figure 5b). Analytical verification of these spectral differences is straightforward due to the simplicity of the system’s dynamics (see section Two nonlinear dynamical systems with subtle dynamical differences, for further details). As we discussed earlier, we expect that Wasserstein-based dynamics similarity analysis is less sensitive to differences between such pairs of dynamical systems as they share identical eigenvalues. However, we expect fastDSA to be sensitive, given that it also considers the flow fields. We empirically validate our hypothesis through simulating these dynamical systems. Given the sensitivity of such systems to the initial condition, we ran the simulations with multiple initial conditions (20 uniformly sampled points near the origin, 80 simulations). We then used Multidimensional Scaling [MDS, cox2008multidimensional] to represent the distance (dissimilarities) derived from fastDSA in a 2-dimensional space. Our results suggest that fastDSA variants correctly distinguish systems A and B (Figure 5c-d), whereas kwDSA struggles to differentiate them (Figure 5f). Furthermore, fastDSA (compared to kwDSA) maintains this separability across a wide range of parameters of our simulation. Given that the separability of these dynamical systems depends on simulation parameters (e.g., the weight of nonlinear terms, $\epsilon$), we used Support Vector Machine [SVM, hearst1998support] with an RBF kernel to separate these two systems across different choices of parameters of our simulations. Our results demonstrate that fastDSA provides superior clustering accuracy compared to kwDSA (Supp. Figure 5). Thus, these results suggest that fastDSA gains efficiency yet remains sensitive to fine changes in underlying dynamics, while alternative routes, based on Wasserstein-based distance, are less sensitive to changes in dynamics.
@@ 3 Discussion
Computationally efficient and accurate methods for comparing dynamics across neural systems are crucial for understanding information processing. Both biological and artificial neural networks process information (encoding, storing, and transforming it) through their complex dynamics [mante2013context, vyas2020computation, luo2025transitions, pagan2025individual, perich2025neural, chaudhuri2019intrinsic, driscoll2024flexible, jensen2024recurrent, ji2025discovering, genkin2025dynamics, tolmachev2025single, sohn2019bayesian, safavi2024signatures, safavi2023brain]. A core method for understanding these processes is dynamic similarity analyses [eisen2024propofol, ostrow2024delay, guilhot2024dynamical, codol2024brain, bolager2024gradient, huang2024measuring, kujur2025transient, xu2025nn, huang2025inputdsa]. FastDSA broadens the applicability of dynamical similarity analysis by achieving high computational efficiency, while maintaining efficacy of the original metric. To realize fastDSA, we incorporate ideas from random matrix theory [plerouRandomMatrixApproach2002, paulRandomMatrixTheory2014, leeRandomMatrixTheory2014, izenmanRandomMatrixTheory, karouiSpectrumEstimationLarge2008], namely, our automated rank estimation through SVHT [gavish2014optimal] and three different variants of accelerated optimizers for the alignment of vector fields. Furthermore, fastDSA reduces computational complexity substantially, while maintaining sensitivity to even fine differences in the dynamics and also being insensitive to purely geometrical changes in the activity. We demonstrate the efficiency, robustness, and sensitivity of all fastDSA variants through a number of numerical experiments [identical or similar to experiments employed by original DSA, ostrow2023beyond]; namely, geometric deformations of a ring attractor network, showing fastDSA methods remain invariant to topology-preserving transformations while being at least $\sim$100x faster than DSA; controlled topological transitions from ring to line attractors, where fastDSA detected changes as effectively as DSA but $\sim$15x faster; and also sensitivity to fine-scale dynamical changes, where fastDSA outperformed previously proposed fast alternatives (kwDSA) in distinguishing systems with identical eigenvalues but different eigenvectors, proving superior sensitivity to subtle differences in underlying dynamics.

### v2

@@ 2.3 Assessing novel optimizers for dynamic alignment comparison
We assessed the performance of the tested methods by computing four summary measures and showed that all were superior to DSA, with some variability among them. We included the following summary measures: (1) the alignment loss, $\|A_{1}-CA_{2}C^{\top}\|_{F}$ (Euclidean score), (2) an angular discrepancy between $A_{1}$ and the aligned $CA_{2}C^{\top}$ to capture qualitative misalignment, (3) the orthogonality residual $\|C^{\top}C-I\|_{F}$, and (4) wall–clock runtime. All hyperparameters, initializations, and stopping criteria were held fixed across methods. The runtime analysis indicated that the regularization-based variant and the Landing algorithm were the most computationally efficient across varying matrix sizes, with the regularization variant being marginally faster (Figure 3a). The Riemannian method was faster than the DSA optimizer for smaller matrices ($n\approx 32$) but exhibited a steeper runtime growth, eventually surpassing DSA’s computational cost for larger values of $n$.
Taken together, our analyses support the Landing-based optimization for solving the alignment problem as the preferred choice in most cases. It matches or improves upon the alignment accuracy of previous dynamic similarity metrics, consistently outperforms both previous methods [57] and the penalty-based and Riemannian variants in wall–clock time, and maintains robust behavior across a wide range of matrix sizes.
2.4 Assessing efficacy and reliability of fastDSA
To demonstrate that the family of fastDSA methods is both efficient and reliable for dynamic similarity analysis, we compare them under diverse scenarios. We apply them to two major sets of simulations, where the underlying dynamics are well characterized a priori (and therefore can serve as a ground truth), and investigate the (in)sensitivity, convergence, computational efficacy, and accuracy of all fastDSA’s variants to different forms of change in governing dynamics.
2.4.1 Invariance under geometric deformations
@@
We first investigated the insensitivity of fastDSA’s variants to purely geometrical transformations (i.e., those that leave the underlying dynamics intact). More specifically, we investigate whether fastDSA maintains invariance under geometric deformations that preserve the attractor topology.
The application of all fastDSA’s methods on geometric deformations that preserve the attractor structure suggests that fastDSA, similar to previous methods [57], maintains invariance under these transformations. We compared the performance and computational efficiency of fastDSA with previous methods [57] based on simulation of a neural system known to represent head direction in the brain, which implements a ring attractor topology embedded in a network of $n$ neurons [92]. We use a conventional network architecture and drive the network using constant input and dynamic noise. Dynamics of the network result in an activity bump moving around the ring attractor at a roughly constant velocity (Figure 4a, left):
@@
Our results suggest that all members of the fastDSA’s family perform equally well as previous methods [DSA and Procrustes analysis 57, in terms of accuracy], while being computationally more efficient. Comparing different variants of fastDSA’s and DSA demonstrates that all methods remain invariant across geometric deformations (Figure 4a, middle; red, green, and purple lines). In contrast, the Procrustes method rapidly diverges with higher values of $\beta$ (Figure 4a, middle; yellow). Furthermore, our results suggest, on average, fastDSA’s computations are at least an order of magnitude faster ($\sim 15\times$) than DSA (Figure 4a, right).
The optimization behavior of the fastDSA family appears to mirror that of DSA, suggesting comparable optimization landscapes and mathematical properties. To evaluate the convergence behavior of different fastDSA methods in comparison to DSA, we conducted a series of numerical experiments. We selected two distinct deformation magnitudes, $\beta=1$ and $\beta=4$ (from Figure 4a, middle). In both cases, we anticipated low dynamical similarity distances, reflecting minor or no topological changes. All fastDSA methods show a decreasing loss (Figure 4b, c left) and lower similarity scores (Figure 4b, c middle), reflecting insensitivity to geometrical deformation. Results indicated that fastDSAmethods consistently converged to stable similarity scores, slightly slower than DSA.
Considering convergence and iteration time of fastDSA  the overall computational efficiency of fastDSA still remains superior to previous methods [57]. To more precisely characterize convergence speed, we analyzed when each algorithm’s loss curve flattened by collapsing run-by-run data into a single mean-loss trajectory per method. For each trajectory, we computed the discrete gradient as the absolute difference between successive loss values, $|\Delta\text{loss}|$, and used a small threshold (0.005) to define when changes become negligible. The stopping iteration was then defined as the first point where $|\Delta\text{loss}|$ dropped below this threshold, or alternatively, the iteration with the minimum gradient if no threshold was crossed. Despite converging slightly later, fastDSA variants were faster overall, requiring substantially less computational time per iteration compared to DSA (see, Figure 4b-c, right, for each iteration time).
@@ 2.4.2 Ability to detect shifts in governing dynamics
On the second step of evaluating fastDSA efficiency and reliability, we investigate the sensitivity of all fastDSA methods to changes in system dynamics; in particular, parametric transformations that alter the topology of the attractors in the dynamical system. To probe the sensitivity of fastDSA we applied a parametrically controlled deformation of the ring topology into a line attractor within the same network model (Figure 4d, left). We implement it by progressively ablating neurons at one end of a recurrent neural network governed by a cosine connectivity kernel. We defined a continuous deformation parameter:
@@
Our results demonstrate that all suggested variants are equally sensitive to change in dynamics [as good as previous method, DSA 57], while being computationally more efficient. Both the family of fastDSA and DSA sharply increased their estimated distance value as $\alpha$ transitioned from less than one (line attractor) to exactly one (ring attractor), confirming their sensitivity to topological changes (Figure 4d, middle; blue, red, green, and purple traces), in contrast to Procrustes analysis (yellow trace). Furthermore, comparing the computation times suggests that fastDSA was approximately 15 times faster than DSA in detecting these topological changes (Figure 4d, right).
@@ 2.4.3 Sensitivity to fine changes in governing dynamics
Lastly, we investigate whether fastDSA is sensitive to fine changes in dynamics and whether this sensitivity is superior to previously suggested alternatives [e.g., Wasserstein-based methods, 36, 65]. A key advantage of fastDSA compared to alternative routes [in terms of computational efficiency, 67, 66, 36] is its sensitivity to minor changes in the dynamics of the system, while maintaining the robustness to noise. The capacity of fastDSA to maintain sensitivity, as elaborated below, stems from its multifaceted approach to computing dynamic similarity. The fastDSA family takes into account both eigenvalues and eigenvectors to compare the governing dynamics; whereas Wasserstein-based approaches [67, 66, 36] focus only on eigenvalue distributions. Thus, as long as the eigenvectors shaping the flow fields are sufficiently structured and distinct, even in the presence of noise, the governing dynamics can be discriminated.
We first introduce one of the key routes for accelerating the estimation of dynamic similarity [discussed in, 36], which is based on a combination of KernelDMD [98, 4], and using the Wasserstein distance as the similarity measure (we refer to it by kwDSA in the following), which brings more computational efficiency to dynamic similarity analysis. We then demonstrate that this alternative approach, kwDSA, is nevertheless less sensitive to changes in dynamics compared to all variants of fastDSA we introduced. Employing kernel-based methods [74, 73] and KernelDMD [25], more specifically, using kernel functions (e.g., Radial Basis Function or RBF) to implicitly map data into a high-dimensional feature space (known as reproducing kernel Hilbert space), considerably enhances computational efficiency, particularly for large-scale or high-dimensional data. This flexibility stems from representing complex nonlinearities without explicitly mapping to the high-dimensional space. By combining KernelDMD with the Wasserstein distance, the similarity between systems can be assessed by comparing the distributions of their spectral components, typically the eigenvalues or singular values of the corresponding Koopman operators [36]. We calculated KernelDMD as
@@
The Wasserstein distance provides a principled way to compare probability distributions and has been used to quantify spectral differences between dynamical systems. A recent study [65] demonstrated that Wasserstein-based metrics can capture dynamical similarity across nonlinear systems, showing both stability and interpretability when comparing spectral measures derived from Koopman or DMD operators. This suggests the Wasserstein distance could be a powerful tool for distribution-based comparisons. However, as we elaborated below, focusing solely on eigenvalue distributions can obscure important geometric differences encoded in the eigenvectors and the underlying flow fields. Thus, it possibly makes the dynamic similarity estimation less sensitive to differences that lie in the configurations of eigenvectors and the associated vector fields. Notably, the distribution of eigenvalues might be very similar across systems, even when the corresponding eigenvectors (which encode the flow directions) differ. To demonstrate the potential shortcomings of kwDSA, we compared the results of kwDSA with fastDSA variants.
@@
Our results suggest that kwDSA is less sensitive to the transition in the regime of the system’s dynamics as the parameter $\alpha$ approaches one (Figure 5a). The advantage of fastDSA over kwDSA is even more pronounced in the presence of noise in the data. To evaluate robustness to noise, we repeated the ring-to-line attractor transition experiment (presented in Figure 4d) while adding Gaussian white noise to the input data (observation noise). We tested four different signal-to-noise ratios (SNRs): 855, 213, 8.3, and 0.6 (corresponding to progressively increasing levels of noise). More specifically, zero-mean white noise was added directly to the neural activity trajectories before applying DMD-based computations. Across all noise levels, fastDSA consistently detected the topological transition (as $\alpha$ approached 1), while kwDSA was less sensitive (Figure 5a; detection capacity degraded after SNR 8.3, second column). These results further highlight fastDSA’s robustness to input variability and its advantage in capturing dynamical changes, even under noisy conditions, compared to fast alternatives, such as kwDSA. Although kwDSA can serve as a reasonable approximation in certain cases [for instance, where comparing only the eigenvalue spectra is sufficient, and we provided an improved implementation of kwDSA in our fastDSA package, 6], it might not be sufficiently sensitive to fine changes in the system’s dynamics; for instance (as discussed in the following) when the dynamics of the systems under comparison only have fine differences in their flow fields.
@@
Lastly, to explicitly demonstrate the mentioned limitation of kwDSA for dynamic similarity estimation (insensitivity to differences in eigenvectors), we also took an analytical approach. We defined two low-dimensional nonlinear systems (system A and system B) with identical eigenvalues but distinct eigenvectors (Figure 5b). Analytical verification of these spectral differences is straightforward due to the simplicity of the system’s dynamics (see section Two nonlinear dynamical systems with subtle dynamical differences, for further details). As we discussed earlier, we expect that Wasserstein-based dynamics similarity analysis is less sensitive to differences between such pairs of dynamical systems as they share identical eigenvalues. However, we expect fastDSA to be sensitive, given that it also considers the flow fields. We empirically validate our hypothesis through simulating these dynamical systems. Given the sensitivity of such systems to the initial condition, we ran the simulations with multiple initial conditions (20 uniformly sampled points near the origin, 80 simulations). We then used Multidimensional Scaling [MDS, 18] to represent the distance (dissimilarities) derived from fastDSA in a 2-dimensional space. Our results suggest that fastDSA variants correctly distinguish systems A and B (Figure 5c-d), whereas kwDSA struggles to differentiate them (Figure 5f). Furthermore, fastDSA (compared to kwDSA) maintains this separability across a wide range of parameters of our simulation. Given that the separability of these dynamical systems depends on simulation parameters (e.g., the weight of nonlinear terms, $\epsilon$), we used Support Vector Machine [SVM, 27] with an RBF kernel to separate these two systems across different choices of parameters of our simulations. Our results demonstrate that fastDSA provides superior clustering accuracy compared to kwDSA (Supp. Figure 5). Thus, these results suggest that fastDSA gains efficiency yet remains sensitive to fine changes in underlying dynamics, while alternative routes, based on Wasserstein-based distance, are less sensitive to changes in dynamics.
@@ 3 Discussion
In this study, we introduce a novel similarity metric, fastDSA, that enables comparing diverse nonlinear dynamical systems at large scales. In particular, the computational efficiency and accuracy of fastDSA, allow comparing dynamics across neural systems with diverse characteristics (e.g,. architectures, input regimes, and learning curricula) to understand their convergent and divergent computational principles [50, 88, 48, 59, 61, 15, 20, 32, 33, 24, 84, 78, 68]. As it was shown in previous studies, a core method for understanding computation in neural systems is dynamic similarity analyses [22, 56, 26, 17, 9, 29, 100, 28] and fastDSA allows these crucial analyses to scale up by achieving high computational efficiency, while maintaining the accuracy of previous metrics [22, 56, 26, 17, 29, 28]. To realize fastDSA, we incorporate ideas from random matrix theory [63, 60, 46, 31, 38], namely, our automated rank estimation through SVHT [23] and three different variants of accelerated optimizers for the alignment of vector fields. Furthermore, fastDSA reduces computational complexity substantially, while maintaining sensitivity to even fine differences in the dynamics and also being insensitive to purely geometrical changes in the activity. We demonstrate the efficiency, robustness, and sensitivity of all fastDSA variants through a number of numerical experiments; namely, geometric deformations of a ring attractor network, showing fastDSA methods remain invariant to topology-preserving transformations while being at least $\sim$150x faster than previous methods [57]; controlled topological transitions from ring to line attractors, where fastDSA detected changes as effectively as previous methods [57] but at least $\sim$15x faster; and also sensitivity to fine-scale dynamical changes, where fastDSA outperformed previously proposed fast alternatives (kwDSA) in distinguishing systems with identical eigenvalues but different eigenvectors, proving superior sensitivity to subtle differences in underlying dynamics. We also drew practical guidance on which alignment optimizer to use when deploying fastDSA at scale. Taking into account the dimensionality of the data, the regularization-based solver and the Landing solver achieved the lowest runtimes, with the regularization solver marginally faster in some regimes. However, the Landing solver exhibited the most stable alignment quality as dimensionality increased. By contrast, both the regularization-based and Riemannian solvers showed degraded reliability beyond small dimensions, and the Riemannian solver additionally exhibited steeper runtime growth that can outweigh its gains at small sizes. Taken together, these findings support the Landing-based optimizer as the default choice for most applications of fastDSA, particularly in higher-dimensional settings and large-sample studies, with the other solvers primarily serving as useful alternatives in small-dimensional regimes or as complementary methods.
The considerable computational efficiency of fastDSA facilitates its application in extensive, large-scale, and statistical analyses, to ultimately better understand neural computations in neural networks. For instance, a recent study demonstrates that dynamic similarity analysis can reveal how computations develop in recurrent neural networks by linking evolving dynamical motifs to learned task structure and performance [26]. It highlights how architectural features and training conditions influence internal dynamics, providing a powerful lens on neural computation. With its improved efficiency, fastDSA enables such analyses at scale, supporting broader studies across architectures, input regimes, and learning curricula.
FastDSA can be particularly helpful to understand computations realized in biological neural systems, as analysis of dynamical similarity provides a crucial tool for validating (or violating) computational models of information processing. For instance, let us consider the example of a Recurrent Neural Network (RNN) trained on a center-out reaching task, against activity recorded from the primary motor cortex (M1) of a monkey performing the same task [17]. Traditional comparison methods, which focus on the geometry (or a static picture) of the high-dimensional neural trajectories, often fail because arbitrary scaling or rotation between the biological and artificial state-spaces can lead to a false conclusion of low similarity, even if the underlying computations are identical. Dynamic similarity analysis can overcome such caveats by comparing the governing dynamics, exploiting the flow field of the system’s dynamics (which, in a way, specifies the instantaneous transformation rule). By establishing whether a single linear map can align the dynamics of the RNN with the dynamics of the M1 neuronal populations, dynamic similarity analysis yields a high similarity score if the RNN is implementing the same computational mechanism as the brain, thus moving validation beyond task performance and to a mechanistic understanding of neural computations. Notably, the computationally efficient (and accurate) similarity metrics are crucial for modern models and datasets, given the growing body of AI-inspired models in neuroscience [e.g., foundation models, 3, 103, 90, 8, 91] and growing complexity and dimensionality of neural data recordings [35, 34, 62, 35].